OpenStudy (anonymous):

THIS QESTION IS IMPOSSIBAL ANYONE WHO CAN ANSWER GETS MEDAL!!!! Owen made 100 sandwiches which she sold for exactly $100. She sold caviar sandwiches for $5.00 each, the bologna sandwiches for $2.00, and the liverwurst sandwiches for 10 cents. How many of each type of sandwich did she make?

OpenStudy (anonymous):

20 caviar sandwiches, and 0 of all other sandwiches.

OpenStudy (anonymous):

nope @OrangeMaster

OpenStudy (anonymous):

Tricky. Set up your equations: C = # of caviar B = # bologna L = # of liverwurst C + B + L = 100 5C + 2B + L/10 = 100 multiply the 2nd equation by 10 and subtract the two 49C + 19B = 900 Now, there are lots of ways to do that, but you need both C and B to be integers. Notice that 900 is divisible by 10. Divide the whole equation by 10 $4.9 C + $1.9 B = $90 Now, we know she need to make exactly $90 selling just caviar and bologna. Each time she sells a sandwich, she'll have to give a dime in change. To get back to an exact dollar figure, she'll need to sell some multiple of 10 sandwiches. So, make the substitution C + B = 10 n Now, solve for C and substitute 49 n - 3 B = 90 Remembering that n and B are both integers, n must be a multiple of 3 (divide that equation by 3 and see). So try, n=3, 6, 9, etc. For n=3, B = 19 For n=6, B = 68 For n=9, B = 117 (which is impossible.) If B = 19, then C = 11 and L = 70. If B = 68, then C = -8 which is impossible. So, the only possible answer is 11 caviar sandwiches, 19 bologna sandwiches and 70 Liverwurst sandwiches.

OpenStudy (anonymous):

What do you mean no?

OpenStudy (anonymous):

i just explained the answer its tricky! @OrangeMaster

OpenStudy (anonymous):

Yeah, that is an answer, but not the only one.

OpenStudy (anonymous):

50 liverwurst, 20 bologna, and 11 caviar.

OpenStudy (anonymous):

no A possible answer can be: 18 Caviar Sandiwiches = $90 4 Bologna Sandwiches = $8 20 Liverwurst Sandwiches = $2 This all equals $100. There can be many more solutions to this problem but this is one of them.

OpenStudy (anonymous):

So, the only possible answer is 11 caviar sandwiches, 19 bologna sandwiches and 70 Liverwurst sandwiches. I know there are more than one solutions, but in the long thing you posted, you said there was only one. And technically my answer would have worked too. You never said that at least on of every sandwich has to be used. You just said 100 sandwiches.

OpenStudy (anonymous):

So, 20 caviar would work.

OpenStudy (anonymous):

very nice!